# SPOT-Trip: Dual-Preference Driven Out-of-Town Trip Recommendation

### v1

@@ Abstract
Out-of-town trip recommendation aims to generate a sequence of Points of Interest (POIs) for users traveling from their hometowns to previously unvisited regions based on personalized itineraries, e.g., origin, destination, and trip duration. Modeling the complex user preferences–which often exhibit a two-fold nature of static and dynamic interests–is critical for effective recommendations. However, the sparsity of out-of-town check-in data presents significant challenges in capturing such user preferences. Meanwhile, existing methods often conflate the static and dynamic preferences, resulting in suboptimal performance. In this paper, we for the first time systematically study the problem of out-of-town trip recommendation. A novel framework SPOT-Trip is proposed to explicitly learns the dual static-dynamic user preferences. Specifically, to handle scarce data, we construct a POI attribute knowledge graph to enrich the semantic modeling of users’ hometown and out-of-town check-ins, enabling the static preference modeling through attribute relation-aware aggregation. Then, we employ neural ordinary differential equations (ODEs) to capture the continuous evolution of latent dynamic user preferences and innovatively combine a temporal point process to describe the instantaneous probability of each preference behavior. Further, a static-dynamic fusion module is proposed to merge the learned static and dynamic user preferences. Extensive experiments on real data offer insight into the effectiveness of the proposed solutions, showing that SPOT-Trip achieves performance improvement by up to 17.01% [^1] The code of SPOT-Trip can be found at https://anonymous.4open.science/r/SPOT-Trip-0607..
@@ 3.2 ODE-Based Dynamic Preference Learning
[…]
@@ 3.3 Static-Dynamic Preference Fusion
$z_{u,n,i}=\phi(\mathbf{W}_{R}[\mathbf{Q}^{o}||\bar{\mathbf{P}}^{o}||\tilde{\mathbf{P}}^{o}]+\mathbf{b}_{R}),$ (10)
@@ 4.1 Experimental Setups
Evaluation Metrics. Following existing studies [14, 23, 34], $F_{1}$ and $PairsF_{1}$, are adopted as the evaluation metrics, where higher values indicate better performance. As the out-of-town trip recommendation focuses on intermediate POIs, we calculate $F_{1}$ and $PairsF_{1}$ without the origin and destination. In addition, to enable more comprehensive comparison, we also report the experiment results with origin and destination, denoted as $Full$-$F_{1}$ and $Full$-$PairsF_{1}$. Please see the corresponding formulas in Appendix C.3.
@@
$F_{1}(\uparrow)$ $PairsF_{1}(\uparrow)$ $Full$-$F_{1}(\uparrow)$ $Full$-$PairsF_{1}(\uparrow)$ $F_{1}(\uparrow)$ $PairsF_{1}(\uparrow)$ $Full$-$F_{1}(\uparrow)$ $Full$-$PairsF_{1}(\uparrow)$
@@ 4.2 Overall Performance
Further, compared to suboptimal baselines equipped with only a single learning module, i.e., Base + KDDC, Base + CNN-ODE and Base + PPROC, SPOT-Trip achieves notable improvements. Specifically, on Foursquare, it improves the $F_{1}$ and $PairsF_{1}$ by 6.67% and 15.96%, respectively, while the corresponding gains are 17.01% and 13.90% on Yelp. Moreover, under full-trip evaluations $Full$-$F_{1}$ and $Full$-$PairsF_{1}$, which inherently include fixed origin and destination and therefore dilute relative gains, SPOT-Trip still consistently outperforms the baselines. These observations demonstrate the superiority of SPOT-Trip due to that it can learn static and dynamic user preferences comprehensively. It is worth noting that Base + PPROC exhibits inferior performance compared to Base + CNN-ODE, potentially because of the complexity involved in predicting spatiotemporal event points, which largely depends on the design of the underlying predictive model.
@@ A.2 Approximate Posterior
where $\mathcal{N}$ is the normal distribution, and $\mathrm{diag}\bigl{(}\psi_{\sigma^{2}}\bigr{)}$ is a diagonal matrix with vector $\psi_{\sigma^{2}}$ on its diagonal. As discussed in Sec. 3.2, $\psi$ is obtained from $\tilde{\mathbf{AGG}}$ with the reparameterization sampling, where we further break up $\psi$ into $[\psi_{\mu},\psi_{\sigma^{2}}]$ as follows:
$\psi_{\mu}=\mathrm{Linear}\bigl{(}\tilde{\mathbf{AGG}}\bigr{)},\psi_{\sigma^{2}}=\exp\Bigl{(}\mathrm{Linear}\bigl{(}\tilde{\mathbf{AGG}}\bigr{)}\Bigr{)},$ (15)
@@ A.3 ELBO
[…]
@@
[…]
@@ C.1 Dataset
#Users #Regions #POIs #Check-ins #Hometown Check-ins #Out-of-town Check-ins #Relations #Entities #Triples
@@ C.3 Evaluation Metrics
As described in Sec. 4.1, our evaluation employs four metrics: $F_{1}$, $PairsF_{1}$, $Full$-$F_{1}$, and $Full$-$PairsF_{1}$. Formally, let $\tau=\{v_{1}^{o},v_{2}^{o},...,v_{N}^{o}\}$ be the generated trips by a method and $\tau^{*}$ be the ground truth, the calculation procedures of these metrics are detailed as follows.
@@
$F_{1}=\frac{2\times Precision\times Recall}{Precision+Recall}$ (17)
@@
Due to the less precise evaluative nature of $Full$-$F_{1}$ and $Full$-$PairsF_{1}$, we report them only in the overall comparison experiments.
@@ C.4 Implementation Details
We re-implemented the baselines and their hyper-parameters based on the details provided in their original papers and publicly available source codes [^4] https://github.com/gcooq/GraphTrip [^5] https://github.com/Mamingqian/MatTrip [^6] https://github.com/Joysmith99/AR-Trip [^7] https://github.com/Yinghui-Liu/KDDC [^8] https://github.com/yakovlev31/pproc-dyn. For the two baselines that require periodic information, Graph-Trip and AR-Trip, we provide additional hour-of-day features as input. For the two baselines that require periodic information, Graph-Trip [14] and AR-Trip [34], we provide additional hour-of-day features as input. In contrast, PPROC [21] relies on both temporal and spatial points to infer preferences at specific spatiotemporal locations. To support this, we additionally design two separate gated recurrent unit (GRU) layers to predict the potential time and location points for future trips. Following [34], we fixed the hidden size of all embeddings to $32$ in all our experiments. The optimizer was uniformly chosen as Adam with an initial learning rate of $0.001$ and L2 regularization with a weight of ${10}^{\textnormal{-}5}$. To avoid overfitting, we adopt the early stop strategy with an 8-epoch patience. In addition, the parameters of our framework and all its variants were consistently set as follows: Consistent with [21], the number of Transformer layers in module ODPL was set to $4$, while both $f(\cdot)$ and $\lambda(\cdot)$ are implemented as $3$-layer MLPs. The latent dimensions of these MLPs were searched from $\{16,32,64,128,256\}$, with $128$ selected as the optimal value based on validation performance. Similarly, we use a differentiable $dopri5$ ODE solver with $rtol$ = $atol$ = $10^{-5}$ from $torchdiffeq$ package [9]. For the static-dynamic preference fusion (Sec. 3.3), the number of Transformer layers was set to $1$, following the configuration in [34]. All Transformer layers employed $4$ attention heads. The hyper-parameter $\sigma_{\tilde{v}^{o}}$ was tuned separately for each dataset, with the optimal value set to $0.6$ for Foursquare and $0.4$ for Yelp. In the optimization stage, the weights of the loss terms $\beta_{1}$, $\beta_{2}$ and $\beta_{3}$ for two datasets were set as $1$.
@@ D.1 Ablation Study
SEEK ($2$[42] 0.0348 0.0082 0.0342 0.0146
SEEK ($4$[42] 0.0376 0.0087 0.0346 0.0154
SEEK ($8$[42] 0.0394 0.0108 0.0361 0.0183
SEEK ($16$[42] 0.0346 0.0073 0.0335 0.0134
@@ D.2 Hyper-parameter Analysis
Figure 8: The effects of the number of hidden states of MLPs in both $f$(·) and $\lambda$(·) on two datasets.
@@
Moreover, we simultaneously vary the latent dimensions of the ODE functions $f$(·) and $\lambda$(·) from $16$ to $256$ to study their joint impact on user dynamic preference modeling in SPOT-Trip. As shown in Fig. 8, we vary the latent dimensions from $16$ to $256$ and evaluate the performance on both Foursquare and Yelp datasets using $F_{1}$ and $PairsF_{1}$ as metrics. On Foursquare, both $F_{1}$ and $PairsF_{1}$ initially decrease at smaller dimensions ($32$ and $64$), but sharply increase at $128$, which yields the best performance. A slight degradation is observed at $256$, likely due to overfitting or increased optimization difficulty. A similar trend is observed on Yelp, where performance significantly improves with larger dimensions and peaks at $128$ before plateauing or dropping slightly at $256$. These findings suggest that the choice of latent dimension also plays a critical role in Neural ODE-based preference modeling. Extremely small dimensions may limit expressiveness, while overly large dimensions can introduce instability. Thus, a latent dimension of $128$ is adopted in our framework.

### v3

@@ Abstract
Out-of-town trip recommendation aims to generate a sequence of Points of Interest (POIs) for users traveling from their hometowns to previously unvisited regions based on personalized itineraries, e.g., origin, destination, and trip duration. Modeling the complex user preferences–which often exhibit a two-fold nature of static and dynamic interests–is critical for effective recommendations. However, the sparsity of out-of-town check-in data presents significant challenges in capturing such user preferences. Meanwhile, existing methods often conflate the static and dynamic preferences, resulting in suboptimal performance. In this paper, we for the first time systematically study the problem of out-of-town trip recommendation. A novel framework SPOT-Trip is proposed to explicitly learns the dual static-dynamic user preferences. Specifically, to handle scarce data, we construct a POI attribute knowledge graph to enrich the semantic modeling of users’ hometown and out-of-town check-ins, enabling the static preference modeling through attribute relation-aware aggregation. Then, we employ neural ordinary differential equations (ODEs) to capture the continuous evolution of latent dynamic user preferences and innovatively combine a temporal point process to describe the instantaneous probability of each preference behavior. Further, a static-dynamic fusion module is proposed to merge the learned static and dynamic user preferences. Extensive experiments on real data offer insight into the effectiveness of the proposed solutions, showing that SPOT-Trip achieves performance improvement by up to 17.01%.
@@ 3.2 ODE-Based Dynamic Preference Learning
$p(\{t_{n}\}^{N}_{n=1}|\tilde{\mathbf{p}}^{o}_{1})=\prod_{n=1}^{N}\lambda(\tilde{\mathbf{p}}^{o}_{n})\mathrm{exp}(-\int\lambda(\tilde{\mathbf{p}}^{o})dt),p(\tilde{\mathbf{v}}^{o}_{n}|\tilde{\mathbf{p}}^{o}_{n},t_{n})=\mathcal{N}(\tilde{\mathbf{p}}^{o}_{n},\sigma^{2}_{\tilde{v}^{o}}I),$ (7)
@@ 3.3 Static-Dynamic Preference Fusion
[…]
@@ 4.1 Experimental Setups
Evaluation Metrics. Following existing studies [14, 23, 34], $F_{1}$ and $PairsF_{1}$, are adopted as the evaluation metrics, where higher values indicate better performance. As the out-of-town trip recommendation focuses on intermediate POIs, we calculate $F_{1}$ and $PairsF_{1}$ without the origin and destination. In addition, to enable more comprehensive comparison, we also report the experiment results with origin and destination, denoted as $Full$-$F_{1}$ and FullFull-$PairsF_{1}$. Please see the corresponding formulas in Appendix C.3.
@@
$F_{1}(\uparrow)$ $PairsF_{1}(\uparrow)$ $F_{1}(\uparrow) FullFull-$$PairsF_{1}(\uparrow)$ $F_{1}(\uparrow)$ $PairsF_{1}(\uparrow)$ $F_{1}(\uparrow) FullFull-$$PairsF_{1}(\uparrow)$
@@ 4.2 Overall Performance
Further, compared to suboptimal baselines equipped with only a single learning module, i.e., Base + KDDC, Base + CNN-ODE and Base + PPROC, SPOT-Trip achieves notable improvements. Specifically, on Foursquare, it improves the $F_{1}$ and $PairsF_{1}$ by 6.67% and 15.96%, respectively, while the corresponding gains are 17.01% and 13.90% on Yelp. Moreover, under full‑trip evaluations $Full$-$F_{1}$ and FullFull-$PairsF_{1}$, which inherently include fixed origin and destination and therefore dilute relative gains, SPOT-Trip still consistently outperforms the baselines. These observations demonstrate the superiority of SPOT-Trip due to that it can learn static and dynamic user preferences comprehensively. It is worth noting that Base + PPROC exhibits inferior performance compared to Base + CNN-ODE, potentially because of the complexity involved in predicting spatiotemporal event points, which largely depends on the design of the underlying predictive model.
@@ A.2 Approximate Posterior
where $\mathcal{N}$ is the normal distribution, and $\mathrm{diag}\bigl(\psi_{\sigma^{2}}\bigr)$ is a diagonal matrix with vector $\psi_{\sigma^{2}}$ on its diagonal. As discussed in Sec. 3.2, $\psi$ is obtained from $\tilde{\mathbf{AGG}}$ with the reparameterization sampling, where we further break up $\psi$ into $\psi_{\mu},\psi_{\sigma^{2}}]$ as follows:
[…]
@@ A.3 ELBO
$=$ $\int q(\tilde{\mathbf{p}}^{o}_{1})\,\ln p\bigl(\tilde{\mathbf{p}}^{o}_{1},\{t_{n},\tilde{\mathbf{v}}^{o}_{n}\}_{n=1}^{N}\bigr)d\tilde{\mathbf{p}}^{o}_{1}$
$+\int q(\tilde{\mathbf{p}}^{o}_{1})\,\ln p(\{{t_{n}}\}_{n-1}^{N}|\tilde{\mathbf{p}}^{o}_{1})d\tilde{\mathbf{p}}^{o}_{1}$
@@
$+\mathbb{E}_{q(\tilde{\mathbf{p}}^{o}_{1})}\sum_{n=1}^{N}\ln\lambda(\tilde{\mathbf{p}}^{o}_{n})-\int\lambda(\tilde{\mathbf{p}}^{o})dt$
@@ C.1 Dataset
#Users #Regions #POIs #Check-ins
#Hometown
Check-ins
#Out-of-town
Check-ins
#Relations #Entities #Triples
@@ C.3 Evaluation Metrics
As described in Sec. 4.1, our evaluation employs four metrics: $F_{1}$, $PairsF_{1}$, FullFull-$F_{1}$, and FullFull-$PairsF_{1}$. Formally, let $\tau=\{v_{1}^{o},v_{2}^{o},...,v_{N}^{o}\}$ be the generated trips by a method and $\tau^{*}$ be the ground truth, the calculation procedures of these metrics are detailed as follows.
@@
$F_{1}=\frac{2\times Precision\times Recall}{Precision+Recall}$ . (17)
@@
Due to the less precise evaluative nature of $Full$-$F_{1}$ and FullFull-$PairsF_{1}$, we report them only in the overall comparison experiments.
@@ C.4 Implementation Details
We re-implemented the baselines and their hyper-parameters based on the details provided in their original papers and publicly available source codes [^3] https://github.com/gcooq/GraphTrip [^4] https://github.com/Mamingqian/MatTrip [^5] https://github.com/Joysmith99/AR-Trip [^6] https://github.com/Yinghui-Liu/KDDC [^7] https://github.com/yakovlev31/pproc-dyn. For the two baselines that require periodic information, Graph-Trip [14] and AR-Trip [34], we provided additional hour-of-day features as input. In contrast, PPROC [21] relies on both temporal and spatial points to infer preferences at specific spatiotemporal locations. To support this, we additionally designed two separate gated recurrent unit (GRU) layers to predict the potential time and location points for future trips. Following [34], we fixed the hidden size of all embeddings to $32$ in all our experiments. The optimizer was uniformly chosen as Adam with an initial learning rate of $0.001$ and L2 regularization with a weight of ${10}^{\textnormal{-}5}$. To avoid overfitting, we adopted the early stop strategy with an 8-epoch patience. In addition, the parameters of our framework and all its variants were consistently set as follows: consistent with [21], the number of Transformer layers in module ODPL was set to $4$, while both $f(\cdot)$ and $\lambda(\cdot)$ are implemented as 33‑layer MLPs. The latent dimensions of these MLPs were searched from $\{16,32,64,128,256\}$, with $128$ selected as the optimal value based on validation performance. Similarly, we used a differentiable $dopri5$ ODE solver with rtolrtol = atolatol = $10^{-5}$ from torchdiffeqtorchdiffeq package [9]. For the static-dynamic preference fusion (Sec. 3.3), the number of Transformer layers was set to $1$, following the configuration in [34]. All Transformer layers employed $4$ attention heads. The hyper-parameter $\sigma_{\tilde{v}^{o}}$ was tuned separately for each dataset, with the optimal value set to $0.6$ for Foursquare and $0.4$ for Yelp. In the optimization stage, the weights of the loss terms $\beta_{1}$, $\beta_{2}$ and $\beta_{3}$ for two datasets were set as $1$.
@@ D.1 Ablation Study
SEEK (22[42] 0.0348 0.0082 0.0342 0.0146
SEEK (44[42] 0.0376 0.0087 0.0346 0.0154
SEEK (88[42] 0.0394 0.0108 0.0361 0.0183
SEEK (1616[42] 0.0346 0.0073 0.0335 0.0134
@@ D.2 Hyper-parameter Analysis
Figure 8: The effects of the number of hidden states of MLPs in both ff(·) and $\lambda$(·) on two datasets.
@@
Moreover, we simultaneously vary the latent dimensions of the ODE functions ff(·) and $\lambda$(·) from $16$ to $256$ to study their joint impact on user dynamic preference modeling in SPOT-Trip. As shown in Fig. 8, we vary the latent dimensions from $16$ to $256$ and evaluate the performance on both Foursquare and Yelp datasets using $F_{1}$ and $PairsF_{1}$ as metrics. On Foursquare, both $F_{1}$ and $PairsF_{1}$ initially decrease at smaller dimensions ($32$ and $64$), but sharply increase at $128$, which yields the best performance. A slight degradation is observed at $256$, likely due to overfitting or increased optimization difficulty. A similar trend is observed on Yelp, where performance significantly improves with larger dimensions and peaks at $128$ before plateauing or dropping slightly at $256$. These findings suggest that the choice of latent dimension also plays a critical role in Neural ODE-based preference modeling. Extremely small dimensions may limit expressiveness, while overly large dimensions can introduce instability. Thus, a latent dimension of $128$ is adopted in our framework.